# 203, B. T. Road, Kolkata-700108, India

### v1

@@ Abstract
A teacher’s knowledge base consists of knowledge of mathematics content, knowledge of student epistemology, and pedagogical knowledge. It has severe implications on the understanding of student’s knowledge of content, and the learning context in general. The necessity to formalize the different content knowledge in approximate senses is recognized in the education research literature. A related problem is that of coherent formalizability. Responsive or smart AI-based software systems do not concern themselves with meaning, and trained ones are replete with their own issues. In the present research, many issues in modeling teachers’ understanding of content are identified, and a two-tier rough set-based model is proposed by the present author. The main advantage of the proposed approach is in its ability to coherently handle vagueness, granularity and multi-modality. An extended example to equational reasoning is used to demonstrate these.
@@
Rough Sets, Education Research, Mathematics Teaching, Equational Reasoning, Mereology, Formalizability Problem, Knowledge Representation, Vague Real Number Systems, Companion Models
@@ 1 Introduction
A number of granular and nongranular semantics of rough sets are known in the literature. Concepts of knowledge can be associated with these from multiple perspectives such as those based on classical rough ideas [33, 30], mereological axiomatic granular perspectives [19, 17, 18], classical granular computing [38], interpretations of modal logic [32, 28], constructive logic [31, 12], concept analysis [39, 8, 18], evidence theory, and machine learning. These ideas of knowledge are not uniformly well-developed across these types of rough sets. Possible applications of these to education research, and specifically mathematical education research is a very domain-specific matter. Many aspects of these are explored by the present author in her earlier papers [21, 20, 27, 16, 23, 24].
@@
In the modern view [1], three components that determine a teacher’s knowledge base are knowledge of mathematics content, knowledge of student epistemology, and pedagogical knowledge. The first is about the breadth and depth of a teacher’s knowledge of the subject. The second concerns understanding of the psychological principles of learning, concept formation, assimilation, and maturation. Teacher’s understanding of teaching in the light of the other components is knowledge of pedagogy. In [10], for example, a very instructive application of the concepts is presented in brief.
@@
Consider the printed contents of a grade six mathematics textbook (of a specific education board). A student’s understanding of the text in relation to the instructional context is distinct from a teacher’s understanding of the pedagogic content (including psychological aspects) and context – the latter, in turn typically differs from a subject expert’s view on the matter (see [1] for a detailed discussion). The categories of students, teachers and experts may be classifiable into a number of finer categories on the basis of subsets of features. The classification of conceptual knowledge as evidenced by the work produced by learners (using any form of expression or in any settings or context) takes many forms. While concept inventories are useful to access over many types of contexts, they do not typically have the breadth to consider atypical reasoning, and related errors. Teachers understanding of the pedagogic content is in relation to all these, and specifically to the meta-level language used for the context. The purpose of this research to propose a framework for this idealized understanding. In actual practice, teachers may themselves have their own misunderstandings, unfounded beliefs and biases, and may lack suitable training for the job. The last aspect is ignored here, and therefore the research is about an idealized understanding. Teachers and students, in a classroom, actually approximate a body of knowledge formed in a distributed cognition perspective in different ways.
@@
The real numbers are taught in certain languages with much interspersed vagueness in middle school. Further, mathematical problem-solving methods, and concepts have vagueness in related discourse. Making use of predicates such as is an approximation of, is definitely a superset of, is definitely a subset of, and is definitely a part of much of the formalization is possible. The harder problem of expressing them through general rough approximations is additionally explored in this research after formalizing the school real numbers typically ignored by formal mathematicians. Work in the area is primarily limited because of the need to systematically confront the varieties of vagueness inherent in it [13, 2]. However, there is much work on pragmatic approaches to the ontology of the real numbers in the context. They serve as sources for building viable models.
@@
In the context of education research in general, and mathematics education research in particular, the very idea of formal language or model is open for some debate. In [24], it is argued that mereology combined with a language of approximations can potentially be used to build higher order formalizations of concepts that go beyond the restrictions envisaged in [13] or in earlier work [2].
@@
Further, a common practice in teaching mathematics is to use everyday language when describing and explaining ideas. Teachers may use phrases such as plug in a value to evaluate a function, and cancel instead of dividing by the common factor. These and the overuse of pronouns can obscure the meaning of the procedures and concepts being used. While learning from errors can be a productive exercise [5], the use of imprecise, loose language, and unidentified errors by teachers can be counterproductive [7, 37]. These and related studies motivate the need to build reasonably formal models to study teachers’ knowledge of content.
The authors of [13], specify the concept of coherent formalizability thus
By coherence, we mean that formalisation of disparate elements hangs together as a meaningful whole. The classroom discourse includes many parts pertaining to definitions, visualisations, representations,
conjectures, exemplification, providing counterexamples, justification, and refutation $\ldots$
@@
Formal models are however not offered in the mentioned paper. This problem is addressed in more detail here through the machinery of partial algebraic systems, mereology, and general rough sets.
@@
To summarize, the following are done in this research (within the domain of pedagogic content knowledge):
@@
    an elaborate justification for the above is argued for through equational reasoning contexts.
@@ 2 Background
Partial operations, and partial algebraic systems pervade this research. For the basics of partial algebras, the reader is referred to [14].
Definition 1
A partial algebra $P$ is a tuple of the form
@@ 2.1 Rough Concept Inventories
Concept inventories are not well-suited for student-centered methods of evaluation because they make use of questions in the multiple-choice query format (MCQ) alone. The reason for choice of incorrect or correct answers can be elicited through additional requirements of explanations [29]. In [21], rough concept inventories are proposed by the present author. These are intended as modification of the methods of a concept inventory for effectively handling vagueness inherent in relatively student centric perspectives. Rough sets are used to represent approximate evaluations.
The methodological aspect is as follows:
  1. 1.
    Select a number of key concepts in a subject or topic;
  2. 2.
    situate them relative to the concepts and granular concepts described in the model in the subsection below (or alternatively situate the concepts relative to a concept map in terms was constructed from and is a part of, and basic well-understood concepts);
  3. 3.
    formulate multiple choice questions that aim to test key aspects of applications of the chosen concepts;
  4. 4.
    each question is required to have at least one correct answer and a number of incorrect answers (distractors) based on student misconceptions or alternative conceptions;
  5. 5.
    require explanation from students for their choice;
  6. 6.
    evaluate explanations relative to the model in terms of concept approximations (or alternatively evaluate explanations relative to concepts that are definitely understood and those that are possibly understood).
@@ 3 The Real Numbers and the Formalizability Problem
The numbers in the braces refer to the place value of operation symbols, and the plus and minus signs have been denoted as $\oplus$ and $\ominus$ respectively. The standard conception of terms in universal algebra is partially split into a number of concepts such as multiplicative terms, binomials, monomials, rational fractions, and fractions in school algebra. Multiplicative terms are also simply referred to as terms, and even these are only implicitly provided with a recursive definition.
The concept of variable symbols and their interpretation over domains is standard in universal algebra. However, the concept of a variable is used in a loose way without a clean specification of the domains in school textbooks.
The predicates $\leq$ and $\geq$ are used in the sense of being numerically less than or being numerically greater than. In the context of word problems, ideas of some objects being bigger or smaller than others often leads to ambiguity and vagueness. This is especially true when objects under consideration have multiple attributes associated. It is also the case that transitive parthood is frequently associated with the $\leq$ relation. Functional non-transitive parthood is apparently not commonly used in mathematics text books at least – though this often happens in the context of activities and descriptions that are very contextual and under-specified. For example, $3$ is the greatest prime less than $5$ and $5$ is the greatest prime less than $7$. However, $3$ is not the greatest prime less than $7$. All this means that the addition of an additional parthood predicate $\mathbf{P}$ to the signature is a good idea. However, the question remains that Where should such predicates be used (since multiple models are intended for the procedure)? Such parthood predicates should be kept in the companion model to avoid complicating the concept of vague reals further. In [27], this aspect is not clarified.
@@ 4 Extended Example: Equational Reasoning
4.1 How and What to Approximate?
Many concepts generated by students are upper approximated by a concept in the real or virtual teacher’s knowledge of content. It is very much possible that relatively unrelated concepts of students are approximated by the same upper approximation. This is not very helpful in practice, as it points to the use of insufficiently fine-grained granules in the process of constructing approximations, and additionally, using an insufficient number of approximation operations. Optimizing over these choices is important, and is illustrated below.
Suppose that a number of students solve the problem Solve $\text{Solve }2x+3=4x+1$. Let six different solutions be as below. Such a scenario may be the case when the problem is offered in a student-centric mode of learning in a larger context. A non-student centric learning algebra context might yield some of the first four responses. Starred steps refer to failed conceptions or attempts at comprehending rules. Each of the solutions can be characterized relative to the concept formation, assimilation, and maturity of the nearest proper solution of the type demonstrated. For example, Solution-5 is a proper version of Solution-1. Part of Solution-1 is correct, and a much greater part of Solution-2 is correct. While predicates such as much greater part of might appear to be a higher order fuzzy predicate, it would be better to characterize them with upper approximations. Writing $(Solution-1)^{uuu}=Solution-5=(Solution-2)^{uu}$ is certainly a meaningful interpretation of a non-idempotent abstract upper approximation. $u$ annihilates negative or insufficient or defective understanding of specific operations or rules, and is therefore a special kind of upper approximation.
@@ 4.1 How and What to Approximate?
$x+3=2x+1$ (*** Cancelling 2 from both sides).
@@
$x=2$ (*** Transposing)
@@
Any comparison of Solution-9 or 6 with the rest requires additional approximate predicates that works across subdomains. Specifically, the comparison of the quality of Solution-6 and Solution-5 can be done through an inventory of common features. This is suggestive of functional parthood predicates [6] such as is a substantial part of or is at least as good as among many others. While the retention of such predicates is useful, it is always better to replace them with ones that are functional (mathematically). More importantly, most teachers would not want to mix up graphical and equational reasoning. Such a desire can be manifested through multiple approximation operators (possibly partial).
@@ 5 The Companion Algebraic Rough Systems
$(\forall x)\mathbf{P}xx$ (PT1)
$(\forall x,b)(\mathbf{P}xb\,\&\,\mathbf{P}bx\longrightarrow x=b)$ (PT2)
@@ 6 Directions
In this research, a formal approach to the problem of reasonably formalizing teachers’ knowledge in middle school mathematics is proposed. The proposed system consists of three major blocks: an improved better suited description of the real numbers as understood in school mathematics, formal systems for reasoning about mathematical assertions that accommodate mereology and vagueness, and psychosocial factors, and a granular extension of the same based on earlier work of the present author. Student-centric concept inventories can additionally be handled more smoothly through the companion models introduced. The problem of coherent formalizability in the education research literature is addressed to an extent.
The purpose of the proposal is to characterize the underlying ontology, improve understanding, and see the impact of many interventions at a formal level. Additionally, it should be usable in focused contexts such as the study of the study of equational reasoning or reasoning about fractions or soft geometry through Geogebra, for a deeper understanding of their potential, and limitations. Future computational models for the contexts, if at all possible, may be assessed through the frameworks introduced. Issues relating to use of multiple approximation operators, ideas of overlap predicates, generalized overlap functions, weak rough implications, rationality, and substantial parthood, will be investigated further in future work.

### v2

@@ Abstract
A teacher’s knowledge base consists of knowledge of mathematics content, knowledge of student epistemology, and pedagogical knowledge. It has severe implications on the understanding of student’s knowledge of content, and the learning context in general. The necessity to formalize the different content knowledge in approximate senses is recognized in the education research literature. A related problem is that of coherent formalizability. Existing responsive or smart AI-based software systems do not concern themselves with meaning, and trained ones are replete with their own issues. In the present research, many issues in modeling teachers’ understanding of content are identified, and a two-tier rough set-based model is proposed by the present author for the purpose of developing software that can aid the varied tasks of a teacher. The main advantage of the proposed approach is in its ability to coherently handle vagueness, granularity and multi-modality. An extended example to equational reasoning is used to demonstrate these. The paper is meant for rough set researchers intending to build logical models or develop meaning-aware AI-software to aid teachers, and education research experts.
@@
Rough Sets, Education Research, Mathematics Teaching, Equational Reasoning, Mereology, Formalizability Problem, Knowledge Representation, Vague Real Number Systems, Companion Models, Rough Logic
@@ 1 Introduction
A number of granular and nongranular semantics of rough sets are known in the literature. Concepts of knowledge can be associated with these from multiple perspectives such as those based on classical rough ideas [32], mereological axiomatic granular perspectives [19, 17, 18], classical granular computing [35], interpretations of modal logic [31, 29], constructive logic [12], concept analysis [36, 8, 18], evidence theory, and machine learning. These ideas of knowledge are not uniformly well-developed across these types of rough sets. Possible applications of these to education research, and specifically mathematical education research is a very domain-specific matter. Many aspects of these are explored by the present author in her earlier papers [20, 27, 16, 22, 23].
@@
In the modern view [1], three components that determine a teacher’s knowledge base are knowledge of mathematics content, knowledge of student epistemology, and pedagogical knowledge. The first is about the breadth and depth of a teacher’s knowledge of the subject. The second concerns understanding of the psychological principles of learning, concept formation, assimilation, and maturation. Teacher’s understanding of teaching in the light of the other components is knowledge of pedagogy. In [10], for example, a very instructive application of the concepts is presented in brief, and it is shown how a teacher’s fixation with some computational procedures for understanding functions can adversely affect conceptual learning outcomes of the class. In the example, the teacher chose not to use the second of the problems suggested (as "it is about chemistry"):
  1. 1.
    A pharmacist is to prepare 15 ml of $2\%$ eye drops for a glaucoma patient. However, she has only $10\%$ and $1\%$ solutions in stock. Can she compound these to fill the prescription?
  2. 2.
    Help her find a convenient way of dispensing different volumes of $2\%$ eye drops for other patients using the same solutions in stock.
Further associations between a teacher’s beliefs about the nature of the subject (such as school mathematics is a fixed set of concepts and procedures that are to be delivered to and remembered by students), and the fixations are deducible. These aspects matter for both the training of, and designing intelligent software aids for teachers (and additionally for the development of curricula and textbooks).
@@
A student’s understanding of a mathematics textbook in relation to the instructional context is distinct from a teacher’s understanding of the pedagogic content (including psychological aspects) and context – the latter, in turn typically differs from a subject expert’s view on the matter (see [1] for a detailed discussion). The categories of students, teachers and experts may be classifiable into a number of finer categories on the basis of subsets of features. The classification of conceptual knowledge as evidenced by the work produced by learners (using any form of expression or in any settings or context) takes many forms. While concept inventories are useful to access over many types of contexts, they do not typically have the breadth to consider atypical reasoning, and related errors. Teachers understanding of the pedagogic content is in relation to all these, and specifically to the meta-level language used for the context. The purpose of this research to propose a framework for this idealized understanding. In actual practice, teachers may themselves have their own misunderstandings, unfounded beliefs and biases, and may lack suitable training for the job. The last aspect is ignored here, and therefore the research is about an idealized understanding. Teachers and students, in a classroom, actually approximate a body of knowledge formed in a distributed cognition perspective [28] in different ways.
@@
The real numbers are taught in certain languages with much interspersed vagueness in middle school. Further, mathematical problem-solving methods, and concepts have vagueness in related discourse. Making use of predicates such as is an approximation of, is a superset of, is a subset of, and is definitely a part of much of the formalization is possible. The harder problem of expressing them through general rough approximations is additionally explored in this research after formalizing the school real numbers typically ignored by formal mathematicians. Work in the area is primarily limited because of the need to systematically confront the varieties of vagueness inherent in it [13, 2]. However, there is much work on pragmatic approaches to the ontology of the real numbers in the context. They serve as sources for building viable models.
@@
In the context of mathematics education research, the very idea of formal language or model is open for some debate. In [23], it is argued that mereology combined with a language of approximations can potentially be used to build higher order formalizations of concepts that go beyond the restrictions envisaged in [13] or in earlier work [2].
@@
Further, a common practice in teaching mathematics is to use everyday language when describing and explaining ideas. Teachers may use phrases such as plug in a value to evaluate a function, and cancel instead of dividing by the common factor. These and the overuse of pronouns can obscure the meaning of the procedures and concepts being used. While learning from errors can be a productive exercise [5], the use of imprecise, loose language, and unidentified errors by teachers can be counterproductive [7]. These and related studies motivate the need to build reasonably formal models to study teachers’ knowledge of content, and more so for the purpose of building intelligent aids for teachers. ML approaches based on numeric simplifications of language features or keywords cannot succeed in the context, and none are attempted in the literature.
@@
The authors of [13], specify the concept of coherent formalizability thus: "By coherence, we mean that formalisation of disparate elements hangs together as a meaningful whole. The classroom discourse includes many parts pertaining to definitions, visualisations, representations, conjectures, exemplification, providing counterexamples, justification, and refutation". Formal models are however not offered in the mentioned paper. This problem is addressed in more detail here through the machinery of partial algebraic systems, mereology, and general rough sets.
@@
To summarize, the following are done in this research (within the domain of pedagogic knowledge):
@@
    a justification of aspects of the above is argued for through equational reasoning contexts.
@@ 2 Background
Partial operations, and partial algebraic systems pervade this research. For the basics of partial algebras, the reader is referred to [14]. A partial algebra $P$ is a tuple of the form
@@ 2.1 Rough Concept Inventories
Concept inventories are not well-suited for student-centered methods of evaluation because they make use of questions in the multiple-choice query format (MCQ) alone. The reason for choice of incorrect or correct answers can be elicited through additional requirements of explanations [30]. In [20], rough concept inventories are proposed by the present author. These are intended as modification of the methods of a concept inventory for effectively handling vagueness inherent in relatively student centric perspectives. Rough sets are used to represent approximate evaluations of explanations. However, the quality of the methodology is bound to depend on that of the underlying language used, and the latter is explored in this research.
@@ 3 The Real Numbers and the Formalizability Problem
The numbers in the braces refer to the place value of operation symbols, and the plus and minus signs have been denoted as $\oplus$ and $\ominus$ respectively. The standard conception of terms in universal algebra is partially split into a number of concepts such as multiplicative terms, binomials, monomials, rational fractions, and fractions in school algebra. Multiplicative terms are also simply referred to as terms, and even these are only implicitly provided with a recursive definition. The concept of variable symbols and their interpretation over domains is standard in universal algebra. However, the concept of a variable is used in a loose way without a clean specification of the domains in school textbooks.
The predicates $\leq$ and $\geq$ are used in the sense of being numerically less than or being numerically greater than. In the context of word problems, ideas of some objects being bigger or smaller than others often leads to ambiguity and vagueness. This is especially true when objects under consideration have multiple attributes associated. It is also the case that transitive parthood is frequently associated with the $\leq$ relation. Functional non-transitive parthood is apparently not commonly used in mathematics text books at least – though this typically happens in the context of activities and descriptions that are very contextual and under-specified. For example, $3$ is the greatest prime less than $5$ and $5$ is the greatest prime less than $7$. However, $3$ is not the greatest prime less than $7$. All this means that the addition of an additional parthood predicate $\mathbf{P}$ to the signature is a good idea. However, the question remains that Where should such predicates be used (since multiple models are intended for the procedure)? Such parthood predicates should be kept in the companion model to avoid complicating the concept of vague reals further. In [27], this aspect is not clarified.
@@ 4 Extended Example: Equational Reasoning
4.1 How to Specify Approximations?
Many concepts generated by students are upper approximated by a concept in the real or virtual teacher’s knowledge of content. It is very much possible that relatively unrelated concepts of students are approximated by the same upper approximation. This is not very helpful in practice, as it points to the use of insufficiently fine-grained granules in the process of constructing approximations, and additionally, using an insufficient number of approximation operations. Optimizing over these choices is important, and is illustrated below. Optionally, approximations can be constructed using the methods of [28] by treating the steps of proofs as granules, adding additional granules, and specifying an up-directed possibly contains relation.
Suppose that a number of students solve the problem Solve $\text{Solve }2x+3=4x+1$. Let six different refined versions of their solutions be as below. Such a scenario may be the case when the problem is offered in a student-centric mode of learning in a larger context. A non-student centric learning algebra context might yield some of the first four responses. Starred steps refer to failed conceptions or attempts at comprehending rules. Each of the solutions can be characterized relative to the concept formation, assimilation, and maturity of the nearest proper solution of the type demonstrated. For example, Solution-5 is a proper version of Solution-1. Part of Solution-1 is correct, and a much greater part of Solution-2 is correct. While predicates such as much greater part of might appear to be a higher order fuzzy predicate, it would be better to characterize them with upper approximations. Writing $(Solution-1)^{uuu}=Solution-5=(Solution-2)^{uu}$ is certainly a meaningful interpretation of a non-idempotent abstract upper approximation. $u$ annihilates negative or insufficient or defective understanding of specific operations or rules, and is therefore a special kind of upper approximation.
@@ 4.1 How to Specify Approximations?
$x+3=2x+1$ (* Cancelling 2 from both sides).
@@
$x=2$ (* Transposing)
The steps of a solution may be seen as concept abstractions, and approximated. Step-2 of Solution-1 and Step-4 of Solution-2 apparently share a similar kind of conceptual error. However, the level of misconception in the former is more than in the latter. Therefore, it is safe to assert that a negation of the latter is part of a negation of the former.
@@
Any comparison of Solution-9 or 6 with the rest requires additional approximate predicates that works across subdomains. Specifically, the comparison of the quality of Solution-6 and Solution-5 can be done through an inventory of common features. This is suggestive of functional parthood predicates [6] such as is a substantial part of [26] or is at least as good as among many others. While the retention of such predicates is useful, it is always better to replace them with ones that are functional (mathematically). More importantly, most teachers would not want to mix up graphical and equational reasoning. Such a desire can be manifested through multiple approximation operators (possibly partial).
@@ 5 The Companion Algebraic Rough Systems
$(\forall x)\mathbf{P}xx;\>(\forall x,b)(\mathbf{P}xb\,\&\,\mathbf{P}bx\longrightarrow x=b)$ (PT1, PT2)
@@ 6 Directions
In this research, a formal approach to the problem of reasonably formalizing teachers’ knowledge in middle school mathematics is proposed. The proposed system consists of three major blocks: an improved better suited description of the real numbers as understood in school mathematics, formal systems for reasoning about mathematical assertions that accommodate mereology and vagueness, and psychosocial factors, and a granular extension of the same based on earlier work of the present author. Student-centric rough concept inventories [20] can additionally be handled more smoothly through the companion models introduced as automatic evaluation of explanations is dependent on the latter. The coherent formalizability issue in the education research literature is thus addressed to an extent. The suggested higher-order formalizations of equational reasoning (including [28]) are not unique, and making these more unique is an important problem.
The purpose of the proposal is to characterize the underlying ontology, improve understanding, and see the impact of many interventions at a formal level. Additionally, it should be usable in focused contexts such as the study of the study of equational reasoning or reasoning about fractions or soft geometry through Geogebra, for a deeper understanding of their potential, and limitations. Future computational models for the contexts, if at all possible, may be assessed through the frameworks introduced. Issues relating to use of multiple approximation operators, ideas of overlap predicates, generalized overlap functions, weak rough implications, rationality, and substantial parthood [26], will be investigated further in future work. The extent of decision-making afforded by the companion systems analogous to those using ortho-pairs [4] is additionally of interest. Additionally, this research motivates higher order approaches in rough sets, that would have been previously glossed over through hybrid layers.